Internal information

vmlinux.org

Prepared by  Document

Joachim Nilsson crash@vmlinux.org L^{A}TEX Template for Technics Package

Approved by  Date

October 12, 2001

L^{A}TEX Template for Technics Package

              Joachim Nilsson crash@vmlinux.org

October 12, 2001

Contents

1 Introduction 2 2 Rationale 2 3 Usage 2 4 Questions & Comments 3

Internal information

vmlinux.org

Prepared by  Document

Joachim Nilsson crash@vmlinux.org L^{A}TEX Template for Technics Package

Approved by  Date

October 12, 2001

1 Introduction

The following two jobs are rolled into one with the modern WYSIWYG^{1} word processor. You type your text, and as you go the text is given a concrete typographical representation on the screen. Supposedly it corresponds closely to what you will see when you send the same document to a printer, although for various reasons it does not always do so. In effect, the text is continuously typeset as you key it in. At first sight this may seem to be a great convenience; on closer inspection it is a curse. There are three aspects to this.

1. The author is distracted from the proper business of composing text, in favor of making typographical choices in relation to which she may have no expertise. “Fiddling with fonts and margins” when she should be focusing on content.

2. The typesetting algorithm employed by WYSIWYG word processor sacrifices quality to the speed required for the setting and resetting of the user’s input in real time. The final product is greatly inferior to that of a real typesetting program.

3. The user of a word processor is under a strong temptation to lose sight of the logical structuration of the text and to conflate this with superficial typographical elements.

2 Rationale

Writing business and technical documents requires the authors full attention to be put on content rather than the above stated issues. L^{A}TEX offers an easy way out of this trap. However, it is criticized for being to cryptic for day to day work by computer illiterates. But do we really want the computer illiterates to dictate the how we write our technical documents?

The obvious answer is an emphatic “No!”. Still, any computer litterate person needs the speed at which todays word processors operate without loosing out to the poor quality those processors performs. To this end the Technics Package was constructed. It provides an advanced typesetting of your technical documents without your being aware of it.

3 Usage

This sample document serves as a template for all you wanting to use the Technics package.

1What You See Is What You Get

Prepared by  Document

Joachim Nilsson crash@vmlinux.org L^{A}TEX Template for Technics Package

Approved by  Date

October 12, 2001

At the very end of this document, i.e., after the \end{document} tag, there are examples of how to construct professional tables, include graphics as well as adding a professionally typeset bibliography to your documents.

4 Questions & Comments

Questions and comments, even improvement patches (though probably not needed :-), can be sent to the author at the below address:

                   Joachim Nilsson <crash@vmlinux.org>